Two resistors of resistance 1 kilo ohm and 2 kilo ohms are in series and a source of 100 volts and negligible resistance is connected to it a voltmeter of 500 ohm is connected in turn across the resistors what are the reading of the voltmeter and how these readings differ from the reading of an ideal voltmeter?
Can someone please give me it's solution

Grade:12

1 Answers

Ketan Prakash
5 years ago
Condintion for ideal voltmeter is R must be infinite.
 
   R eq.  = 3500 ohms.
  Apply ohm's law , V=iR
               i=100/3500 A = 1/35 A
Current flowing in the series is i=1/35 A
Since combination in series so all current is same everywhere.
        In voltmeter when it is not ideal having Resistance 500 ohm , V=1/35 x 500 V
                                       V=14.29v
Voltmeter when it is ideal means R equals to infinite ,i.e. ,V=iR = tends to infinite.